# Scaling Linear Attention with Sparse State Expansion

### v1

@@ Abstract
The Transformer architecture, despite its widespread success, struggles with long-context scenarios due to quadratic computation and linear memory growth. While various linear attention variants mitigate these efficiency constraints by compressing context into fixed-size states, they often degrade performance in tasks such as in-context retrieval and reasoning. To address this limitation and achieve more effective context compression, we propose two key innovations. First, we introduce a row-sparse update formulation for linear attention by conceptualizing state updating as information classification. This enables sparse state updates via softmax-based top-$k$ hard classification, thereby extending receptive fields and reducing inter-class interference. Second, we present Sparse State Expansion (SSE) within the sparse framework, which expands the contextual state into multiple partitions, effectively decoupling parameter size from state capacity while maintaining the sparse classification paradigm. Our design, supported by efficient parallelized implementations, yields effective classification and discriminative state representations. We extensively validate SSE in both pure linear and hybrid (SSE-H) architectures across language modeling, in-context retrieval, and mathematical reasoning benchmarks. SSE demonstrates strong retrieval performance and scales favorably with state size. Moreover, after reinforcement learning (RL) training, our 2B SSE-H model achieves state-of-the-art mathematical reasoning performance among small reasoning models, scoring 64.7 on AIME24 and 51.3 on AIME25, significantly outperforming similarly sized open-source Transformers. These results highlight SSE as a promising and efficient architecture for long-context modeling.
@@ 1 Introduction
The Transformer architecture [48] has achieved remarkable success in various sequence modeling tasks, leveraging its attention mechanism for expressive modeling and high parallelism. However, Transformers inherently struggle with processing long sequences due to quadratic computational complexity and linearly growing memory usage of key-value (KV) caches during inference.
@@
Motivated by the pursuit of more effective context compression, we propose two key innovations. First, we introduce a row-sparse update formulation for linear attention by conceptualizing contextual state updating as information classification. Specifically, we interpret the key vector mapping $\mathbf{k}_{t}=f(\mathbf{x}_{t},\mathbf{W}_{k})$ as a classification function and employ a softmax-based top-$k$ hard classification strategy. At each step, this formulation selectively updates only state rows corresponding to predicted classes. Both theoretical analysis and synthetic experiments validate that our approach achieves larger receptive fields and reduces inter-class interference. Second, to mitigate the memory capacity constraints caused by limited state sizes, we extend the contextual state within the row-sparse update framework and propose Sparse State Expansion (SSE). SSE expands the state into $N$ partitions with shared attention parameters and utilizes a learnable bias for partition selection, followed by a softmax-based selection of state rows within partitions. This design effectively decouples parameter size from state capacity while maintaining the sparse classification paradigm, and empirical results indicate that the states of SSE achieves lower inter-row similarity and higher singular value entropy, which enables more discriminative state representations. We also propose efficient parallelized implementations of SSE, employing masking and the varlen technique optimized for various training contexts.
We extensively validate SSE under both linear and hybrid architectures (denoted as SSE-H), across three core capabilities: language modeling, in-context retrieval, and mathematical reasoning. SSE demonstrates strong language modeling performance among advanced linear models [56, 55, 13], while SSE-H outperforms both hybrid and Transformer baselines. On long-context retrieval tasks, SSE consistently improves upon other linear attention models, and its hybrid variant significantly narrows the gap with softmax attention Transformers. For reasoning, 2B SSE-H model achieves scores of 64.7 on AIME24 and 51.3 on AIME25 with advanced RL training [42, 58]—exceeding the best reported results from similarly sized open-source Transformers (Figure 1. Left). Moreover, SSE exhibits strong scalability with respect to state capacity (Figure 1. Right), and supports efficient conversion from pretrained Transformers at different scales, demonstrating its flexibility across model architectures and training regimes.
@@
  • We extensively validate SSE and the hybrid SSE-H across diverse training stages and benchmark tasks. Our 2B hybrid model achieves state-of-the-art reasoning performance among small reasoning models, scoring 64.7 on AIME24 and 51.3 on AIME25—exceeding previous open-source softmax attention Transformers of comparable size.
@@ 2 Preliminary
Softmax Attention. Softmax attention [48] stores the key-value vectors derived from historical tokens in a KV cache ($\mathbf{K}_{t}\in\mathcal{R}^{t\times d}$,$\mathbf{V}_{t}\in\mathcal{R}^{t\times d}$). As a new token $\mathbf{x}_{t}$ arrives, the corresponding vectors $\mathbf{k}_{t},\mathbf{v}_{t}$ are appended to this cache. Attention is then computed by evaluating the interaction between the current token’s query $\mathbf{q}_{t}$ and the KV cache:
$\mathbf{S}_{t}$ =$=\{\mathbf{K}_{t},\mathbf{V}_{t}\},$ (1)
@@ 4 SSE: State Expansion under the Row-Sparse Framework
Figure 4: Comparison between vanilla linear attention and SSE. SSE expands the state into $N$ partitions under the proposed row-sparse framework. All partitions share attention parameters; sparse row selection is determined by top-$k$ partition selection based on a learnable bias, followed by softmax over selected rows.
@@ 4.1 Sparse State Expansion via Information Classification
Within the information classification framework, a straightforward strategy involves expanding the number of classes by a factor of $N$, resulting in $N\times c$ total state rows organized into $N$ partitions. For each input token $\mathbf{x}_{t}$, we utilize the top-$k$-then-softmax mechanism to select a subset of rows from the chosen partitions. Building on our prior analysis, this design is driven by two key objectives: (1) scaling the state size independently of parameter count, and (2) preserving the sparse classification paradigm by explicitly performing information classification through $\operatorname{softmax}(\operatorname{top\text{-}}k(\mathbf{x}_{t}\mathbf{W}_{k}))$. Accordingly, our core design consists of the following components:
Shared attention parameters across partitions. We use a single set of projection weights across all partitions. This design choice is equivalent to performing segmented clustering of input information along the sequence dimension, thereby enabling sparse token-level interactions.
Learnable bias for hard partition selection, followed by soft state row selection via softmax. Top-$k$: We first compute a bias vector $\mathbf{e}_{t}\in\mathcal{R}^{N}$ via a linear transformation of the input, which is then used to select the top-$k$ partitions. Then softmax: To preserve the unified classification behavior, we apply softmax over all state rows corresponding to the selected partitions. Since all partitions share the same projection matrix $\mathbf{W}_{k}$, applying softmax over the $k\times c$ selected state rows is mathematically equivalent to applying softmax independently within each of the $k$ partitions and scaling the result by a factor of $\frac{1}{k}$. To ensure the bias parameters are trainable, we directly incorporate the bias term into the softmax input. Given the translation invariance of softmax, this bias can be embedded into the classification function without altering the output distribution, whose scale is governed by softmax. As a result, the bias serves exclusively for top-$k$ selection while maintaining consistent classification behavior. For a selected partition $i$, the output key vector is computed as: $\mathbf{k}_{t}^{i}=\frac{1}{k}\operatorname{softmax}(\mathbf{x}_{t}\mathbf{W}_{k}+\mathbf{e}_{t}^{i})$.
@@
$\mathbf{e}_{t}$ $=\mathbf{x}_{t}\mathbf{W}_{e},\ \mathcal{T}=\{i\ |\ \mathbf{e}_{t}^{i}\in\operatorname{top\text{-}}k(\mathbf{e}_{t})\}$ (9)
@@
$\mathbf{k}_{t}$ $=\frac{1}{k}\operatorname{softmax}(\mathbf{x}_{t}\mathbf{W}_{k}+\mathbf{e}_{t}^{i}),\quad\text{for}\ i\in\mathcal{T},$ (11)
$\mathbf{S}_{t}^{i}$ $=\mathbf{\Lambda}_{t}\mathbf{S}_{t-1}^{i}+\mathbf{k}_{t}^{\top}\mathbf{v}_{t},\quad\text{for}\ i\in\mathcal{T}.$ (12)
@@
Beyond state updating, we can also perform state reading selectively based on information similarity. Within the state expansion framework, a simple yet effective strategy is to restrict reading to the top-$k^{\prime}$ partitions, utilizing the same bias-based selection with $\mathbf{e}_{t}$: $\mathbf{o}_{t}=\sum_{i\in\mathcal{T^{\prime}}}\mathbf{q}_{t}\mathbf{S}_{t}^{i}$, where $\mathcal{T^{\prime}}=\{i\ |\ \mathbf{e}_{t}^{i}\in\operatorname{top-k}^{\prime}(\mathbf{e}_{t})\}$. By default, and unless otherwise specified, we set $\mathcal{T^{\prime}}=\mathcal{T}$. Leveraging parameter sharing and the sparsity induced by hard classification, SSE enables scaling up the state size of linear attention while keeping both computational and parameter overhead nearly constant.
Remarks. Local interactions between consecutive tokens serve as an effective prior for language modeling [15, 34, 3], which we leverage to stabilize the sparse component’s training. To this end, we incorporate an always-selected partition and employ a LoRA strategy to maintain a nearly constant parameter count [^7] LoRA is applied only to the QK projections, while V projections are shared across all partitions, including the always-selected one.. Furthermore, the partition selection mechanism (i.e., the computation of $\mathbf{e}_{t}$) offers flexibility for incorporating diverse inductive biases into the model design. For instance, this allows integrating positional and historical information beyond the current context-only formulation $\mathbf{e}_{t}=\mathbf{x}_{t}\mathbf{W}_{e}$, a direction for future work.
@@ 5 Experiments
Wiki.$\downarrow$ PIQA Hella. Wino. ARC-e ARC-c SIQA Avg. FDA SWDE SQuAD Avg.
@@ 5.1 Language Modeling and Retrieval
Language Modeling. We conducte small-scale pretraining to compare different attention mechanisms, using 600M and 2B models trained on 15B and 100B tokens, respectively. Subsequently, we evaluate these models on zero-shot commonsense reasoning performance. As showed in Table 1, SSE-n4 demonstrates strong results among advanced linear attention models, outperforming Transformers and matching MoM despite using fewer parameters (600M vs. 800M). Furthermore, in a hybrid setting, SSE-H also surpasses both GLA-H and Transformers at the 2B scale, reaching state-of-the-art performance.
@@ 5.2 Reasoning Ability
As shown in Table 5, our 2B SSE-H-n4k1 achieves substantial gains over similarly sized softmax Transformers—improving AIME24 accuracy from 48.3 to 64.7 (+16.4), and AIME25 from 36.8 to 51.3 (+14.5). Its performance also approaches that of much larger models like DeepSeek-R1-Distill-Qwen-7B [20], underscoring the strength of hybrid models for high-level reasoning.
@@
SSE-H-n4k1-2B (Ours) 64.7 51.3 92.8 86.0 92.0
5.3 Scalability of State Size
How does the recall performance of SSE scale with state size? We evaluate this by conducting experiments on 2B models trained on 100B tokens, varying both the number of partitions ($n$) and the top-$k$ hard selection size ($k$), and measuring the corresponding changes in recall performance. The results are presented in Figure 7. Under the sparse update framework, state expansion in SSE exhibits promising scalability without increasing parameter count:
@@ 5.3 Scalability of State Size
While demonstrating promising scalability, increasing the sparsity ratio and partition number incurs additional computational costs and inference latency. Consequently, we employed the $n4k1$ setting for the majority of our empirical analysis, particularly for the hybrid architecture. This indicates that SSE’s performance holds significant room for further improvement.

### v2

@@ Abstract
The Transformer architecture, despite its widespread success, struggles with long-context scenarios due to quadratic computation and linear memory growth. While various linear attention variants mitigate these efficiency constraints by compressing context into fixed-size states, they often degrade performance in tasks such as in-context retrieval and reasoning. To address this limitation and achieve more effective context compression, we propose two key innovations. First, we introduce a row-sparse update formulation for linear attention by conceptualizing state updating as information classification. This enables sparse state updates via softmax-based top-$k$ hard classification, thereby extending receptive fields and reducing inter-class interference. Second, we present Sparse State Expansion (SSE) within the sparse framework, which expands the contextual state into multiple partitions, effectively decoupling parameter size from state capacity while maintaining the sparse classification paradigm. Supported by efficient parallelized implementations, our design achieves effective classification and highly discriminative state representations. We extensively validate SSE in both pure linear and hybrid (SSE-H) architectures across language modeling, in-context retrieval, and mathematical reasoning benchmarks. SSE demonstrates strong retrieval performance and scales favorably with state size. Moreover, after reinforcement learning (RL) training, our 2B SSE-H model achieves state-of-the-art mathematical reasoning performance among small reasoning models, scoring 64.5 on AIME24 and 50.2 on AIME25, significantly outperforming similarly sized open-source Transformers. These results highlight SSE as a promising and efficient architecture for long-context modeling.
@@ 1 Introduction
The Transformer architecture [48] has achieved remarkable success in various sequence modeling tasks, leveraging its attention mechanism for expressive modeling and high parallelism. However, it inherently struggles with processing long sequences due to quadratic computational complexity and linearly growing memory usage of key-value (KV) caches during inference.
@@
Motivated by the pursuit of more effective context compression, we propose two key innovations. First, we introduce a row-sparse update formulation for linear attention by conceptualizing contextual state updating as information classification. Specifically, we interpret the key vector mapping $\mathbf{k}_{t}=f(\mathbf{x}_{t},\mathbf{W}_{k})$ as a classification function and employ a softmax-based top-$k$ hard classification strategy. At each step, this formulation selectively updates only state rows corresponding to predicted classes. Both theoretical analysis and synthetic experiments validate that our approach achieves larger receptive fields and reduces inter-class interference. Second, to mitigate the memory capacity constraints caused by limited state sizes, we extend the contextual state within the row-sparse update framework and propose Sparse State Expansion (SSE). SSE expands the state into $N$ partitions with shared attention parameters and utilizes a write-read gate for partition selection, followed by a softmax-based selection of state rows within partitions. This design effectively decouples parameter size from state capacity while maintaining the sparse classification paradigm, and empirical results indicate that the states of SSE achieves lower inter-row similarity and higher singular value entropy, which enables more discriminative state representations. We also propose efficient parallelized implementations of SSE, employing masking and the varlen technique optimized for various training contexts.
We extensively validate SSE under both linear and hybrid architectures (denoted as SSE-H), across three core capabilities: language modeling, in-context retrieval, and mathematical reasoning. SSE demonstrates strong language modeling performance among advanced linear models [56, 55, 13], while SSE-H outperforms both hybrid and Transformer baselines. On long-context retrieval tasks, SSE consistently improves upon other linear attention models, and its hybrid variant significantly narrows the gap with softmax attention Transformers. For reasoning, 2B SSE-H model achieves scores of 64.5 on AIME24 and 50.2 on AIME25 with advanced RL training [42, 58]—exceeding the best reported results from similarly sized open-source Transformers (Figure 1. Left). Moreover, SSE exhibits strong scalability with respect to state capacity (Figure 1. Right), and supports efficient conversion from pretrained Transformers at different scales, demonstrating its flexibility across model architectures and training regimes.
@@
  • We extensively validate SSE and the hybrid SSE-H across diverse training stages and benchmarks. Our 2B hybrid model achieves state-of-the-art reasoning performance among small reasoning models, scoring 64.5 on AIME24 and 50.2 on AIME25—exceeding previous open-source softmax attention Transformers of comparable size.
@@ 2 Preliminary
Softmax Attention. Softmax attention [48] stores the key-value vectors derived from historical tokens in a KV cache ($\mathbf{K}_{t}\in\mathcal{R}^{t\times d}$$\mathbf{V}_{t}\in\mathcal{R}^{t\times d}$). As a new token $\mathbf{x}_{t}$ arrives, the corresponding vectors $\mathbf{k}_{t},\mathbf{v}_{t}$ are appended to this cache. Attention is then computed by evaluating the interaction between the current token’s query $\mathbf{q}_{t}$ and the KV cache:
$\mathbf{S}_{t}$ $=\{\mathbf{K}_{t},\mathbf{V}_{t}\},$ (1)
@@ 4.1 Sparse State Expansion via Information Classification
Figure 4: Comparison between vanilla linear attention and SSE. SSE expands the state into $N$ partitions within the row-sparse update framework, where a classification function assigns information to specific state rows. All partitions share attention parameters. Sparse row selection follows two steps: (1) top-$k$ partition selection based on a write-read gate (blue indicates selected partitions; green marks an always-selected partition for training stability), and (2) row selection within the chosen partitions via softmax over key vectors.
Within the information classification framework, a straightforward strategy involves expanding the number of classes by a factor of $N$, resulting in $N\times c$ total state rows organized into $N$ partitions. For each input token $\mathbf{x}_{t}$, we utilize the top-$k$-then-softmax mechanism to select a subset of rows from the chosen partitions. Building on our prior analysis, this design is driven by two key objectives: (1) scaling the state size independently of parameter count, and (2) preserving the sparse classification paradigm by explicitly performing information classification. For efficiency, we avoid applying top-$k$ directly across all state rows (Equation 7). Instead, we first perform hard top-$k$ selection over partitions, followed by soft row selection within the chosen partitions using softmax. Accordingly, our core design consists of the following components:
Shared attention parameters across partitions. We use a single set of projection weights (QKV) across all partitions. This design choice is equivalent to performing segmented clustering of input information along the sequence dimension, thereby enabling sparse token-level interactions.
Write-read gate for hard partition selection, followed by soft row selection via softmax. A gate vector $\mathbf{e}_{t}\in\mathcal{R}^{N}$ is first used to select the top-$k$ partitions, and within these partitions $\operatorname{softmax}(\mathbf{k}_{t})$ further selects the state rows. Only the selected partitions are updated, and ablation results show that the softmax step yields over 40% row sparsity. Furthermore, the gate is applied to both state input (KV) and output (Q), thereby governing both information writing and reading.
@@
$\mathbf{e}_{t}$ $=\operatorname{softmax}(\mathbf{x}_{t}\mathbf{W}_{e}),$ (9)
$\mathcal{T}$ $=\{i\ |\ \mathbf{e}_{t}^{i}\in\operatorname{top\text{-}}k(\mathbf{e}_{t})\},$ (10)
@@
$\mathbf{k}_{t}$ $=\operatorname{softmax}(\mathbf{x}_{t}\mathbf{W}_{k}),$ (12)
[…]
$\mathbf{o}_{t}$ $=\sum_{i\in\mathcal{T}}\mathbf{e}_{t}^{i}\cdot\mathbf{q}_{t}\mathbf{S}_{t}^{i}.$ (14)
(1) We compute a gate vector $\mathbf{e}_{t}\in\mathcal{R}^{N}$ through a linear transformation of the input followed by softmax (Equation 9), and select the top-$k$ partitions based on $\mathbf{e}_{t}$ (Equation 10). (2) To preserve the unified classification behavior, we apply softmax over all state rows of the selected partitions. Since all partitions share the same projection matrix $\mathbf{W}_{k}$, applying softmax over the $k\times c$ selected rows is equivalent to applying it independently within each partition (Equation 12); the scaling factor can be absorbed into the parameters and thus omitted. (3) To ensure the gate remains trainable, we incorporate $\mathbf{e}_{t}$ into both state input (KV) and output (Q) (Equations 1314), thereby controlling both writing and reading and enabling more targeted optimization. (4) For a selected partition $i$, the gated key vector is $\mathbf{k}_{t}^{i}=\mathbf{e}_{t}^{i}\cdot\operatorname{softmax}(\mathbf{x}_{t}\mathbf{W}_{k})$ (Equation 13). By the normalization property of softmax, the gate can be embedded into the classification function while preserving consistent behavior, with classification intensity $\sum_{i}\sum_{d}(\mathbf{k}_{t}^{i})_{d}=1$. Only the selected partitions $\mathbf{S}_{t}^{i}$ are updated. (5) Beyond state updating, we also perform selective reading based on information similarity. Within the write–read gating framework, we restrict reading to the top-$k^{\prime}$ partitions using the same gate $\mathbf{e}_{t}$ (Equation 14): $\mathbf{o}_{t}=\sum_{i\in\mathcal{T^{\prime}}}\mathbf{e}_{t}^{i}\cdot\mathbf{q}_{t}\mathbf{S}_{t}^{i}$, where by default $\mathcal{T^{\prime}}=\mathcal{T}$.
@@
Local interactions between consecutive tokens serve as an effective prior for language modeling [15, 34, 3], which we leverage to stabilize the sparse component’s training. To this end, we incorporate an always-selected partition and employ a LoRA strategy to maintain a nearly constant parameter count [^7] LoRA is applied only to the QK projections, while V projections are shared across all partitions, including the always-selected one.. Leveraging parameter sharing and the sparsity induced by hard classification, SSE enables scaling up the state size of linear attention while keeping both computational and parameter overhead nearly constant.
@@ 5 Experiments
[…]
@@ 5.1 Language Modeling and Retrieval
Language Modeling. We conduct small-scale pretraining to compare different attention mechanisms, using 600M and 2B models trained on 15B and 100B tokens, respectively. Subsequently, we evaluate these models on zero-shot commonsense reasoning performance. As shown in Table 1, SSE-n4 demonstrates strong results among advanced linear attention models, outperforming Transformers and matching MoM despite using fewer parameters (600M vs. 800M). Furthermore, in a hybrid setting, SSE-H also surpasses both GLA-H and Transformers at the 2B scale, reaching state-of-the-art performance.
@@ 5.2 Reasoning Ability
As shown in Table 5, our 2B SSE-H-n4k1 achieves substantial gains over similarly sized open-source softmax Transformers—improving AIME24 accuracy from 48.3 to 64.5 (+16.2), and AIME25 from 36.8 to 50.2 (+13.4). Its performance also approaches that of much larger models like DeepSeek-R1-Distill-Qwen-7B [20], underscoring the strength of hybrid models for high-level reasoning.
@@
SSE-H-n4k1-2B (Ours) 64.5 50.2 92.1 85.7 91.4
5.3 Ablation Study
Effectiveness of State Size Scaling in SSE Recall. We evaluate this by conducting experiments on 2B models trained on 20B tokens, varying both the number of partitions ($n$) and the top-$k$ hard selection size ($k$), and measuring the corresponding changes in recall performance. The results are presented in Figure 7. Under the sparse update framework, state expansion in SSE exhibits promising scalability without increasing parameter count:
While demonstrating promising scalability, increasing the sparsity ratio and partition number incurs additional computational costs and inference latency. Consequently, we employed the $n4k1$ setting for the majority of our empirical analysis, particularly for the hybrid architecture. This indicates that SSE’s performance still exhibits substantial potential for further improvement.
@@ 5.3 Ablation Study
Figure 8: Row sparsity ratio of state updating. The row sparsity ratio is measured by the proportion of channels in $\mathbf{k}_{t}$ with values below $1\text{e-}5$.
Effectiveness of Softmax-based Row Selection. We conduct an ablation study on 600M models trained with 15B tokens to evaluate the effectiveness of the proposed softmax feature map for row-sparse updates under state expansion. As Table 6 shows, using softmax as the feature map for $\mathbf{k}_{t}$ yields significantly higher recall compared to the commonly used SiLU map. This is consistent with our earlier analysis based on information classification.
We further examine the sparsity of $\mathbf{k}_{t}$ vectors in 2B GLA and SSE-n4k1 models, using $1\text{e-}5$ as the threshold. As shown in Figure 8, although SiLU has a value range near zero, the trained GLA model does not exhibit any sparsity. In contrast, SSE with the softmax classifier induces sparsity in the selected partitions, with ratios of 5% and 42%. Interestingly, the model learns to write densely to the shared partition while maintaining highly sparse updates in the non-shared partitions.
Table 6: Ablation study on the effect of softmax-based row selection.
Models Params. Tokens FDA SWDE SQuAD Avg.
SSE-n4k1-k.silu 600M 15B 20.87 26.91 24.90 24.23
SSE-n4k1-k.softmax 600M 15B 33.67 33.30 26.51 31.16 (+6.93)
Effectiveness of Write-Read Gating. Within SSE’s hard partition selection framework, we introduce a write-read gate that operates on both the state input and output (i.e., KV and Q) simultaneously. We conduct a comprehensive ablation study on 2B models trained with 100B tokens, comparing four variants: (1) no gating, (2) input-only write gate, (3) output-only read gate, and (4) our proposed write-read gate. As Table 7 illustrates, the write-read gate consistently outperforms all alternatives on both recall and commonsense reasoning tasks, demonstrating the effectiveness of partition selection design in SSE.
Table 7: Ablation study on write-read gating. Models: SSE with write–read gating (applied to both input and output) compared against three variants: no gate, write-only gate, and read-only gate. ✓indicates the gate is applied, and ✗ indicates it is not. Metrics: CSR-Avg. is the average of commonsense reasoning tasks.
[…]
SSE-n4k1 2B 100B 18.44 64.07 65.17 40.65 56.63 53.63
no gate 2B 100B 18.63 51.91 63.91 39.81 51.87 (-4.76) 53.01 (-0.62)
write gate 2B 100B 18.65 54.54 61.12 37.10 50.92 (-5.71) 52.82 (-0.81)
read gate 2B 100B 18.62 57.26 64.45 40.18 53.96 (-2.67) 52.91 (-0.72)
Effectiveness of Shared Attention Parameters. SSE shares attention parameters (QKV and gate projection) across partitions, while applying LoRA to decouple QK projections for the always-selected partition. For comparison, we train a variant where each of the $N+1$ partitions has its own attention parameters, denoted as "wo/ shared-params". As shown in Table 8, despite nearly doubling the non-embedding parameters, this variant shows lower recall and only marginal gains on commonsense reasoning. These results support the key assumption of SSE: overcoming the limitations of linear attention depends on state capacity rather than parameter count. Effective parameter sharing not only enables the model to learn segmented clustering of input information, but may also facilitate more stable optimization.
Table 8: Ablation study on shared attention parameters. Models: SSE shares attention parameters across partitions, while "wo/ shared-params" denotes a variant with separate parameters for each partition. Metrics: CSR-Avg. is the average of commonsense reasoning tasks.
[…]
SSE-n4k1 600M (300M) 15B 35.52 33.67 33.30 26.51 31.16 42.91
wo/ shared-params 890M (580M) 15B 34.94 16.52 36.45 24.87 25.95 (-5.21) 42.99 (+0.08)